OpenStudy (anonymous):

Can someone explain how to convert between s,v,a,and delta, omega, and alpha?

OpenStudy (roadjester):

If I'm understanding this correctly then s=displacement v=ds/dt=velocity a=dv/dt=acceleration I don't know about delta \[\omega=angular\space\space velocity=v/r\] \[\alpha=angular\space\space acceleration=a/r\]

OpenStudy (roadjester):

If these variables are consistent, then I would guess that delta is your angular position but I'm used to seeing theta, not delta